OpenStudy (anonymous):

A rectangle initially has width 4 meters and length 8 meters and is expanding so that the area increases at a rate of 9 square meters per hour. If the width increases by 20 centimeters per hour how quickly does the length increase initially?

OpenStudy (anonymous):

Use the area of a rectangle equation: \[A = lw\] find the first derivative (rate of change of area) equation \[A\prime = l(w)\prime + (l)\prime w\] Solve for \[(l)\prime \] using the information

OpenStudy (anonymous):

i got 50

OpenStudy (anonymous):

did you convert the 20 cm per hour to 0.20 m per hour?

OpenStudy (anonymous):

i believe i did, what did you come up with?

OpenStudy (anonymous):

I got 1.85

OpenStudy (anonymous):

1.85 m per hour

OpenStudy (anonymous):

Did you use \[A \prime = 9 m/hr, w = 4 m, l = 8m, w\prime = 0.20 m/hr\]

OpenStudy (anonymous):

Yea, i meant the final answer. I got 50

OpenStudy (anonymous):

look at your algebra steps that can not be correct it is to large

OpenStudy (anonymous):

ok so with 1.85, you did what after that?

OpenStudy (anonymous):

That is what I get for the final answer after using the above numbers

OpenStudy (anonymous):

@PhysicsGuru I understand why that should be the correct answer, but I don't understand why it doesn't work out when you try to take that answer and go back and figure out the area after x hours. For instance, after 1 hour, at that rate, the length would be 9.85 and the width would be 4.2, which should be 32 + 9, but it's not quite. 9.85 x 4.2 = 41.37.

OpenStudy (anonymous):

where are you getting 32 + 9?

OpenStudy (anonymous):

The area at the beginning (4 meters x 8 meters = 32 square meters) plus the area after one hour, growing at a rate of 9 square meters per hour. 32 + 9 = area after one hour.

OpenStudy (anonymous):

@PhysicsGuru can you just give me the answer and I can understand how it was derived from their